OpenStudy (amistre64):

What does the scalar result of dotting 2 vectors signify? I know a 0 means they are perps, but what is the scalar product actually mean?

OpenStudy (amistre64):

<1,1>.<0,2sqrt(2)> = 0 + 2sqrt(2) = 2sqrt(2). That doesnt make sense to me a whole lot.

OpenStudy (amistre64):

that just seems to give the magnitude of the bottom vector

OpenStudy (anonymous):

<1,1>.<0,2sqrt(2)> = 0 + 2sqrt(2) = 2sqrt(2). how?

OpenStudy (nowhereman):

It gives the cosine of the angle between them. This is the way you can get angles in any vector space with an inner product.

OpenStudy (amistre64):

<1 , 1 > <0,2sqrt(2)> ------------ 0 + 2sqrt(2) = 2sqrt(2)

OpenStudy (amistre64):

yes, it gives the adjacent part of the cosine angle which is just the magnitude of the bottom vector right? the product of the magnitudes of the top and bottom vectors account for the 'hypotenuse' then

OpenStudy (anonymous):

<1 , 1 > <0,2sqrt(2)> ------------ 0 + 2sqrt(2) = 2sqrt(2). iam not geting it amistre

OpenStudy (amistre64):

see; it is how you multiply vectors; or at least one way to do it; the other being the cross product

myininaya (myininaya):

cross product can be found two ways if i remember correctly finding the determinant of a matrix or some kind of trig thing right?

OpenStudy (amistre64):

determinate of 2 vectors is usual

OpenStudy (amistre64):

the other way is just to equate the zero dot product of 2 vectors inthe plane and equat them

OpenStudy (amistre64):

v<x,y> n<a,b> ------- vx.a + vy.b = 0 u<x,y> n<a,b> -------- ux.a +uy.b = 0 equate them and solve the system of equations :)

OpenStudy (amistre64):

geterminate is mor emechanical tho

OpenStudy (amistre64):

lol... read thru the typos

OpenStudy (amistre64):

i now the scalar projection projects the top onto the bottom at |top|cos(a)

OpenStudy (amistre64):

i just dont understand if there is another way to use the scalar product other than associating it to the cos(a) between the 2 vectors :)

myininaya (myininaya):

example: find u*v if u=2i+3j and v=-2k | i j k | = i|3 0 | -j|2 0| +k|2 3| =i(-2(3)-0)-j(-2-0)+k(0-0)=-6i+2j | 2 3 0 | |0 -2| |0 -2| |0 0| | 0 0 -2|

OpenStudy (amistre64):

correct; just keep in mind to do +-+ :) that middle - is to account for being lazy i think lol

myininaya (myininaya):

=-6j+4k

myininaya (myininaya):

oops =-6i+4j

myininaya (myininaya):

i forgot about that othr two up there

OpenStudy (amistre64):

2a+3b+0c = 0 0a+0b-2c = 0 lol... thats would be hard to determine that way ; cross product is better in that case; and in most cases prolly

myininaya (myininaya):

yes that would be hard lol (or impossible) i dont see how we could do it that way